solve the following equation for the given value of x:
y= cos x; for x = 4 pie

Found 2 solutions by rfer, Alan3354:
Answer by rfer(12730)

y=cos x
y=cos(4*3.14)
y=cos 12.56
y=0.976

Answer by Alan3354(31528)

y= cos x; for x = 4 pie
----
y = cos(x) = cos(4pi) = 1
It's pi, the Greek letter, not pie, the dessert.